# Local problems in trees across a wide range of distributed models

### v1

@@ Abstract
  • any model that does not violate physical causality [Gavoille, Kosowski, Markiewicz, DICS 2009],
@@ 1 Introduction
While the randomized online-$\operatorname{\mathsf{LOCAL}}$ model is in general much stronger than the deterministic $\operatorname{\mathsf{LOCAL}}$ model, in this work we show that in trees, for many families of local problems these two models (and hence all models in between) are asymptotically equally strong. Figure 3 summarizes the relations that we have thanks to this work; for comparison, Figure 2 shows the state of the art before this work.
@@ 1.2 Prior Work on LCL Problems
  • In 2-dimensional grids, 3-coloring has polynomial locality in e.g. randomized $\operatorname{\mathsf{SLOCAL}}$ and non-signaling models, while it can be solved with logarithmic locality in deterministic online-$\operatorname{\mathsf{LOCAL}}$ [2].
@@ 1.3 Contributions
In all these settings, it is known that any LCL problem falls in one of the following complexity classes in the deterministic $\operatorname{\mathsf{LOCAL}}$ model: the locality is $O(1)$, $\Theta(\operatorname{log}^{*}n)$, $\Theta(\operatorname{log}n)$, or $\Theta(n^{1/k})$ for some $k=1,2,3,\dots$ [15, 11, 9, 4, 3]. Furthermore, in the case of rooted regular trees, we can (relatively efficiently) also decide in which of these classes any given LCL problem belongs to [3].
In this work we show that in many of these complexity classes, deterministic $\operatorname{\mathsf{LOCAL}}$ and randomized online-$\operatorname{\mathsf{LOCAL}}$ are asymptotically equally strong. Our main contributions are:
@@
By prior work, the relation between deterministic $\operatorname{\mathsf{LOCAL}}$ and randomized online-$\operatorname{\mathsf{LOCAL}}$ was well-understood in the $o(\operatorname{log}\operatorname{log}\operatorname{log}n)$ region for rooted (not necessarily regular) trees [1]; see Figure 2 in the appendix for the state of the art before this work. Putting together prior results and new results, the landscape shown in Figure 3 emerges.
@@ 2 Key Ideas, Technical Overview, and Comparison with Prior Work
We keep the discussion at a high level but invite the interested reader to consult the formal definitions on Section 3 as needed.
@@
If we could prove a similar statement for every gap result for LCLs in trees, we would then have a similar implication for each complexity class: the complexities across all models in Figure 1 would be the same for every LCL problem in trees. Alas, such a statement cannot be true in full generality (as we discussed above, there are some known separations between the models), but in this work we take major steps in many cases where that seems to be the case. To understand how we achieve this, it is useful to make a distinction between two flavors of prior work: classification-style and speedup-style arguments; we will make use of both in this work.
@@ 2.1 Classification Arguments for Regular Trees
Note that we do not need to change the third step; as soon as we establish property $P$, the preexisting deterministic $\operatorname{\mathsf{LOCAL}}$ algorithm kicks in. This is, in essence what we do in Sections 4, 6 and 7: we take the prior classification from [3] for rooted and unrooted regular trees and show that it can be extended to cover the entire range of models all the way from deterministic $\operatorname{\mathsf{LOCAL}}$ to randomized online-$\operatorname{\mathsf{LOCAL}}$.
@@
One key property that we make use of is so-called certificates of $O(\operatorname{log}^{*}n)$ solvability, introduced in [4]. In Section 4 we show that the existence of an $o(\operatorname{log}n)$-locality randomized online-$\operatorname{\mathsf{LOCAL}}$ algorithm implies that the LCL problem has a certificate of $O(\operatorname{log}^{*}n)$ solvability, which (as the name suggests) implies that the problem can be solved with $O(\operatorname{log}^{*}n)$ locality in both the $\operatorname{\mathsf{CONGEST}}$ and $\operatorname{\mathsf{LOCAL}}$ models. This then extends to an $O(1)$ upper bound in $\operatorname{\mathsf{SLOCAL}}$ model [13] and in deterministic online-$\operatorname{\mathsf{LOCAL}}$ [2], which then directly implies the same upper bound also for randomized online-$\operatorname{\mathsf{LOCAL}}$ model. This way we can establish the following result:
@@
Let $\Pi$ be an LCL problem on rooted regular trees. If $\Pi$ can be solved with $o(\operatorname{log}n)$ locality in randomized online-$\operatorname{\mathsf{LOCAL}}$, then it can be solved in $\operatorname{\mathsf{LOCAL}}$ with $O(\operatorname{log}^{*}n)$ locality. Consequently, $\Pi$ can be solved with $O(1)$ locality in $\operatorname{\mathsf{SLOCAL}}$ (and thus also with the same locality in online-$\operatorname{\mathsf{LOCAL}}$, even deterministically).
@@
Another key property that we make use of is the depth of an LCL (see Section 5): to every LCL problem $\Pi$ on regular trees there is an associated quantity $d_{\Pi}$ called its depth that depends only on the description of $\Pi$ and which allows us to classify its complexity in the deterministic $\operatorname{\mathsf{LOCAL}}$ model [3]. We show that depth also captures the complexity in randomized online-$\operatorname{\mathsf{LOCAL}}$. In Section 6 we first analyze the case of rooted regular trees and show the following:
@@
The high-level strategy is as follows: assuming we have an algorithm with locality $o(n^{1/k})$, if it does not fail, then its existence would imply $d_{\Pi}>k$, which is a contradiction. In Section 7 we prove an analogous statement for unrooted regular trees:
@@
  • If $0<d_{\Pi}<\infty$, then $\Pi$ has locality $\Theta(n^{1/k})$ in both $\operatorname{\mathsf{LOCAL}}$ and randomized online-$\operatorname{\mathsf{LOCAL}}$.
  • If $d_{\Pi}=\infty$, then $\Pi$ can be solved in $\operatorname{\mathsf{CONGEST}}$ (and hence also in online-$\operatorname{\mathsf{LOCAL}}$, even deterministically) with locality $O(\operatorname{log}n)$.
@@
    We extend the classification all the way to randomized online-$\operatorname{\mathsf{LOCAL}}$, while [2] only discusses deterministic online-$\operatorname{\mathsf{LOCAL}}$. While this may at first seem like a technicality, Figure 1 highlights the importance of this distinction: randomized online-$\operatorname{\mathsf{LOCAL}}$ captures all models in this diagram. This includes in particular the quantum-$\operatorname{\mathsf{LOCAL}}$ model and the non-signaling model. Note it is currently open if deterministic online-$\operatorname{\mathsf{LOCAL}}$ captures either of these models.
@@
      We can extend our work to unrooted trees, whereas the results in [2] only apply to rooted trees.
@@ 2.2 Speedup Arguments for General Trees
Here a key challenge is that we need to explicitly change models; while in general deterministic $\operatorname{\mathsf{LOCAL}}$ is not strong enough to simulate randomized online-$\operatorname{\mathsf{LOCAL}}$, we show that in this case such a simulation is possible. This is, in essence, what we do in Section 8: we start with the argument from [5] and show that sublinear-locality randomized online-$\operatorname{\mathsf{LOCAL}}$ algorithms not only admit speedup inside randomized online-$\operatorname{\mathsf{LOCAL}}$, but also a simulation in deterministic $\operatorname{\mathsf{LOCAL}}$. Formally, we prove:
@@
Suppose $\Pi$ is an LCL problem that can be solved with $o(n)$ locality in the online-$\operatorname{\mathsf{LOCAL}}$ model (with or without randomness) in unrooted trees. Then $\Pi$ can be solved in unrooted trees in online-$\operatorname{\mathsf{LOCAL}}$ with $O(\sqrt{n})$ locality (with or without randomness, respectively).
@@ 3 Preliminaries
We denote the set of natural numbers, including zero, by $\mathbb{N}$. For positive numbers integers, excluding zero, we use $\mathbb{N}_{+}$. For a set or multiset $X$ and $k\in\mathbb{N}_{+}$, we write ${\bigl{(}\!{\binom{X}{k}}\!\bigr{)}}$ for the set of multisets over $X$ of cardinality $k$.
In this work we are concerned exclusively with simple graphs. Unless stated otherwise, $n$ always denotes the number of nodes in the graph. An algorithm solving some graph problem is said to succeed with high probability if, for all graphs except finitely many, it fails with probability at most $1/n$.
@@ 3.1 Locally Checkable Labeling Problems
In this section we define locally checkable labeling (LCL) problems as well as some of their key analytical concepts. We first recall the most general definition due to Naor and Stockmeyer [19].
@@
For a graph $G=(V,E)$ whose vertices are labeled according to $\Sigma_{\mathrm{in}}$, a (node) labeling of a graph $G=(V,E)$ is a solution to $\Pi$ if it labels every node $v\in V$ with a label from $\Sigma_{\mathrm{out}}$ such that the $r$-neighborhood of $v$ in $G$ is identical to some graph of $\mathcal{C}$ (when we place $v$ at the center of the respective graph in $\mathcal{C}$). Every node for which this requirement holds is said to be correctly labeled. If the nodes of $G$ are labeled with a solution, then $G$ itself is correctly labeled.
@@
Under this formalism, leaves may be labeled arbitrarily. When $\Pi$ is clear from the context, we refer to the elements of ${\bigl{(}\!{\binom{\Sigma}{\Delta}}\!\bigr{)}}$ as node configurations.
Meanwhile, in the case of unrooted trees, we work with LCL problems where the labels are placed on (instead of nodes) and are subject to node and edge constraints.
@@
In the same vein as the preceding definition, leaves are unconstrained and may be labeled arbitrarily. As before, if such an LCL problem $\Pi$ on regular unrooted trees is clear from the context, we refer to the elements of ${\bigl{(}\!{\binom{\Sigma}{\Delta}}\!\bigr{)}}$ as node configurations and to those of ${\bigl{(}\!{\binom{\Sigma}{2}}\!\bigr{)}}$ as edge configurations.
@@ 3.2 Models of Distributed Computing
The deterministic $\operatorname{\mathsf{LOCAL}}$ model of distributed computing runs on a graph $G=(V,E)$ resembling a computer network, with each node $v\in V$ representing a computer, and each edge representing a connection channel. Each node is labeled with a unique identifier. All nodes run the same distributed algorithm in parallel. Initially, a node is only aware of its own identifier and degree. Computation proceeds in synchronous rounds, and in each round a node can send and receive a message to and from each neighbor and update its state. Eventually each node must stop and announce its local output (its part of the solution, e.g. in graph coloring its own color). The running time, round complexity, or locality of the algorithm is the (worst-case) number of rounds $T(n)$ until the algorithm stops in any $n$-node graph.
@@ 4 Sub-logarithmic Gap for LCLs in Rooted Regular Trees
Throughout this section, we consider an LCL problem $\Pi=(\Delta,\Sigma,\mathcal{V})$ and a randomized online-$\operatorname{\mathsf{LOCAL}}$ algorithm $\mathcal{A}$ that solves $\Pi$ with locality $T(n)=o(\operatorname{log}n)$ with high probability. We start by constructing a family of input instances. We then argue that $\mathcal{A}$ solving these instances must produce a canonical labeling regardless of the randomness. Finally, we show how we can use this canonical labeling to construct the coprime certificate for $O(\operatorname{log}^{*}n)$ solvability.
@@ 4.1 Construction of Input Instances
We construct the family of input instances as follows:
@@
Observe that trees constructed in this way have $n=(\lvert\Sigma\rvert+1)\cdot\Delta^{d+1}\cdot\frac{\Delta^{2d+1}-1}{\Delta-1}$ nodes. Moreover, choosing $(b,u,C)$ uniquely fixes the construction. Let $\mathcal{P}$ be the set of choices for $(b,u,C)$.
@@ 4.2 Randomness of the Algorithm
We now analyze the failure probability of $\mathcal{A}$ to show that there exists a fixed labeling $\overline{\sigma*}$ for middle nodes $M$ such that the labeling is still completable for each input instance $(b,u,C)$. Let $\mathcal{F}$ denote the event that algorithm $\mathcal{A}$ fails to solve the problem. As, by assumption, $\mathcal{A}$ succeeds with high probability, we must have that $\Pr_{S,\mathcal{A}^{+}(b,u,C)}(\mathcal{F})\leq\frac{1}{n}$ for all choices of $(b,u,C)\in\mathcal{P}$. We can now pick a labeling of the middle nodes that minimizes the average failure probability:
@@
$\frac{\lvert\mathcal{P}\rvert}{n}\geq\sum_{(b,u,C)\in\mathcal{P}}\Pr_{\mathcal{A}^{+}(b,u,C)}(\mathcal{F}\mid(b,u,C),S=\overline{\sigma*})$
@@
Hence for each instance $(b,u,C)$, the algorithm fails with a probability strictly less than 1. In other words, it must succeed with non-zero probability, and hence a correct labeling must also exist.

### v2

@@ Abstract
  • any model that does not violate physical causality [Gavoille, Kosowski, Markiewicz, DISC 2009],
@@ 1 Introduction
While the randomized online-$\operatorname{\mathsf{LOCAL}}$ model is in general much stronger than the deterministic $\operatorname{\mathsf{LOCAL}}$ model, in this work we show that in trees, for many families of local problems these two models (and hence all models in between) are asymptotically equally strong. Figure 2 summarizes the relations that we have thanks to this work; for comparison, see Figure 3 for the state of the art before this work.
@@ 1.2 Prior Work on LCL Problems
  • In 2-dimensional grids, 3-coloring has polynomial locality in e.g. randomized $\operatorname{\mathsf{SLOCAL}}$ and non-signaling models, while it can be solved with logarithmic locality in deterministic online-$\operatorname{\mathsf{LOCAL}}$ [2, 12].
@@ 1.3 Contributions
In all these settings, it is known that the locality of any LCL problem in deterministic $\operatorname{\mathsf{LOCAL}}$ falls in one of the following classes: $O(1)$, $\Theta(\operatorname{log}^{*}n)$, $\Theta(\operatorname{log}n)$, or $\Theta(n^{1/k})$ for some $k=1,2,3,\dots$ [16, 11, 9, 4, 3]. Furthermore, in the case of rooted regular trees, we can (relatively efficiently) also decide which of these classes any given LCL problem belongs to [3].
Our main contribution is showing that, in many of these complexity classes, deterministic $\operatorname{\mathsf{LOCAL}}$ and randomized online-$\operatorname{\mathsf{LOCAL}}$ are asymptotically equally strong:
@@
By prior work, the relation between deterministic $\operatorname{\mathsf{LOCAL}}$ and randomized online-$\operatorname{\mathsf{LOCAL}}$ was well-understood in the $o(\operatorname{log}\operatorname{log}\operatorname{log}n)$ region for rooted (not necessarily regular) trees [1]; see Figure 3 for the state of the art before this work. Putting together prior results and new results, the landscape shown in Figure 2 emerges.
@@ 2 Key Ideas, Technical Overview, and Comparison with Prior Work
We keep the discussion at a high level but invite the interested reader to consult the formal definitions in Section 3 as needed.
@@
If we could prove a similar statement for every gap result for LCLs in trees, we would then have a similar implication for each complexity class: the complexities across all models in Figure 1 would be the same for every LCL problem in trees. Alas, such a statement cannot be true in full generality (as we discussed above, there are some known separations between the models), but in this work we take major steps in many cases where that seems to hold. To understand how we achieve this, it is useful to make a distinction between two flavors of prior work: classification- and speedup-style arguments; we will make use of both techniques.
@@ 2.1 Classification Arguments for Regular Trees
Note that we do not need to change the third step; as soon as we establish property $P$, the pre-existing deterministic $\operatorname{\mathsf{LOCAL}}$ algorithm kicks in. This is, in essence, what we do in Sections 4, 6 and 7: we take the prior classification from [3] for rooted and unrooted regular trees and show that it can be extended to cover the entire range of models all the way from deterministic $\operatorname{\mathsf{LOCAL}}$ to randomized online-$\operatorname{\mathsf{LOCAL}}$.
@@
One key property that we make use of are certificates of $O(\operatorname{log}^{*}n)$ solvability for rooted regular trees, introduced in [4]. In Section 4, we show that the existence of an $o(\operatorname{log}n)$-locality randomized online-$\operatorname{\mathsf{LOCAL}}$ algorithm implies that the LCL problem has such a certificate; in turn, as the name suggests, the certificate of solvability implies that the problem can be solved with $O(\operatorname{log}^{*}n)$ locality in both $\operatorname{\mathsf{CONGEST}}$ and $\operatorname{\mathsf{LOCAL}}$. This then extends to an $O(1)$ upper bound in $\operatorname{\mathsf{SLOCAL}}$ [14] and in deterministic online-$\operatorname{\mathsf{LOCAL}}$ [2], which then directly implies the same upper bound also for randomized online-$\operatorname{\mathsf{LOCAL}}$ model. Hence we obtain the following:
@@
Let $\Pi$ be an LCL problem on rooted regular trees. If $\Pi$ can be solved with $o(\operatorname{log}n)$ locality in randomized online-$\operatorname{\mathsf{LOCAL}}$, then it can be solved in $\operatorname{\mathsf{LOCAL}}$ with $O(\operatorname{log}^{*}n)$ locality. Consequently, $\Pi$ can be solved with $O(1)$ locality in $\operatorname{\mathsf{SLOCAL}}$, and thus also with the same locality in online-$\operatorname{\mathsf{LOCAL}}$, even deterministically.
@@
Another key property that we make use of is the depth of an LCL (see Section 5): to every LCL problem $\Pi$ on regular trees there is an associated quantity $d_{\Pi}$ called its depth that depends only on the description of $\Pi$ and which allows us to classify its complexity in the deterministic $\operatorname{\mathsf{LOCAL}}$ model [3]. We show that depth also captures the complexity in randomized online-$\operatorname{\mathsf{LOCAL}}$. In Section 6, analyze the case of rooted regular trees and show the following:
@@
The high-level strategy is as follows: Assume we have an algorithm with locality $o(n^{1/k})$. We construct a large experiment graph and run the algorithm on that. If the algorithm fails, then it contradicts the assumption that we have such algorithm. If the algorithm succeeds, then it also produces a certificate showing that $d_{\Pi}>k$, which again is a contradiction. In Section 7, we prove an analogous statement for unrooted regular trees:
@@
  • If $0<d_{\Pi}<\infty$, then $\Pi$ has locality $\Theta(n^{1/d_{\Pi}})$ in both deterministic $\operatorname{\mathsf{LOCAL}}$ and randomized online-$\operatorname{\mathsf{LOCAL}}$.
  • If $d_{\Pi}=\infty$, then $\Pi$ can be solved in $\operatorname{\mathsf{CONGEST}}$– and hence also in online-$\operatorname{\mathsf{LOCAL}}$, even deterministically – with locality $O(\operatorname{log}n)$.
@@
    We extend the classification all the way to randomized online-$\operatorname{\mathsf{LOCAL}}$, while [2] only discusses deterministic online-$\operatorname{\mathsf{LOCAL}}$. While this may at first seem like a technicality, Figure 1 highlights the importance of this distinction: randomized online-$\operatorname{\mathsf{LOCAL}}$ captures all models in this diagram. This includes in particular the quantum-$\operatorname{\mathsf{LOCAL}}$ model and the non-signaling model. Note that it is currently open if deterministic online-$\operatorname{\mathsf{LOCAL}}$ captures either of these models.
@@
      Our work extends to unrooted trees; the results in [2] only apply to rooted trees.
@@ 2.2 Speedup Arguments for General Trees
Here a key challenge is that we need to explicitly change models; while in general deterministic $\operatorname{\mathsf{LOCAL}}$ is not strong enough to simulate randomized online-$\operatorname{\mathsf{LOCAL}}$, we show that in this case such a simulation is possible. This is, in essence, what we do in Section 8: starting with the argument from [5], we show that sublinear-locality randomized online-$\operatorname{\mathsf{LOCAL}}$ algorithms not only admit speedup inside randomized online-$\operatorname{\mathsf{LOCAL}}$, but also a simulation in deterministic $\operatorname{\mathsf{LOCAL}}$. Formally, we prove:
@@
Suppose $\Pi$ is an LCL problem that can be solved with $o(n)$ locality in online-$\operatorname{\mathsf{LOCAL}}$ (with or without randomness) in unrooted trees. Then $\Pi$ can be solved in unrooted trees in $\operatorname{\mathsf{LOCAL}}$ with $O(\sqrt{n})$ locality (with or without randomness, respectively).
@@ 3 Preliminaries
[…]
All graphs are simple. Unless stated otherwise, $n$ always denotes the number of nodes in the graph. An algorithm solving some graph problem is said to succeed with high probability if, for all graphs except finitely many, it fails with probability at most $1/n$.
@@ 3.1 Locally Checkable Labeling Problems
In this section we define locally checkable labeling (LCL) problems. We first recall the most general definition due to Naor and Stockmeyer [20]:
@@
For a graph $G=(V,E)$ whose vertices are labeled according to $\Sigma_{\mathrm{in}}$, a (node) labeling of a graph $G=(V,E)$ is a solution to $\Pi$ if it labels every node $v\in V$ with a label from $\Sigma_{\mathrm{out}}$ such that the $r$-neighborhood of $v$ in $G$ is identical to some graph of $\mathcal{C}$ (when we place $v$ at the center of the respective graph in $\mathcal{C}$). Every node for which this holds is said to be correctly labeled. If all nodes of $G$ are labeled with a solution, then $G$ itself is correctly labeled.
@@
[…]
Meanwhile, in the case of unrooted trees, we work with LCL problems where the labels are placed on half-edges (instead of nodes) and are subject to node and edge constraints.
@@
[…]
@@ 3.2 Models of Distributed Computing
The deterministic $\operatorname{\mathsf{LOCAL}}$ model of distributed computing runs on a graph $G=(V,E)$ where each node $v\in V$ represents a computer and each edge a connection channel. Each node is labeled with a unique identifier. All nodes run the same distributed algorithm in parallel. Initially, a node is only aware of its own identifier and degree. Computation proceeds in synchronous rounds, and in each round a node can send and receive a message to and from each neighbor and update its state. Eventually each node must stop and announce its local output (its part of the solution, e.g. in graph coloring its own color). The running time, round complexity, or locality of the algorithm is the (worst-case) number of rounds $T(n)$ until the algorithm stops in any $n$-node graph.
@@ 4 Sub-logarithmic Gap for LCLs in Rooted Regular Trees
Throughout this section, we consider an LCL problem $\Pi=(\Delta,\Sigma,\mathcal{V})$ and a randomized online-$\operatorname{\mathsf{LOCAL}}$ algorithm $\mathcal{A}$ that solves $\Pi$ with locality $T(n)=o(\operatorname{log}n)$ with high probability. We start by constructing a family of input instances and then argue that $\mathcal{A}$ solving these instances produces a canonical labeling regardless of the randomness. Finally, we show how this canonical labeling yields the coprime certificate for $O(\operatorname{log}^{*}n)$ solvability.
@@ 4.1 Construction of Input Instances
To construct the family of input instances:
@@
Note that trees constructed in this way have $n=(\lvert\Sigma\rvert+1)\cdot\Delta^{d+1}\cdot\frac{\Delta^{2d+1}-1}{\Delta-1}$ nodes. Moreover, choosing $(b,u,C)$ uniquely fixes the construction. Let $\mathcal{P}$ be the set of choices for $(b,u,C)$.
@@ 4.2 Randomness of the Algorithm
We now analyze the failure probability of $\mathcal{A}$ to show that there exists a fixed labeling $\overline{\sigma*}$ for middle nodes $M$ such that the labeling is still completable for each input instance $(b,u,C)$. Let $\mathcal{F}$ denote the event that algorithm $\mathcal{A}$ fails to solve the problem. By assumption, $\mathcal{A}$ succeeds with high probability; hence $\Pr_{S,\mathcal{A}^{+}(b,u,C)}(\mathcal{F})\leq\frac{1}{n}$ for all choices of $(b,u,C)\in\mathcal{P}$. We can now pick a labeling of the middle nodes that minimizes the average failure probability:
@@
$\frac{\lvert\mathcal{P}\rvert}{n}\geq\sum_{(b,u,C)\in\mathcal{P}}\Pr_{\mathcal{A}^{+}(b,u,C)}(\mathcal{F}\mid(b,u,C),S=\overline{\sigma*}).$
@@
Hence for each instance $(b,u,C)$, the algorithm fails with probability strictly less than 1. Thus it succeeds with non-zero probability, and hence a correct labeling must exist.